For the following scores, find the (a) mean, (b) median, (c) sum of squared deviations, (d) variance, and (e) standard deviation. 6,-1,3,-3,4
Answer #1

Solution Given that,

x x2
6 36
-1 1
3 9
-3 9
4 16
\sumx = 9 \sumx2 = 71

a ) Mean =  \sumx / n

= 6 -1+3-3+4 / 5

= 9 / 5

= 1.8

Mean = 1.8

b ) Median

Observations in the ascending order are :
-3,-1,3,4,6

Here, n = 5 is odd.

M = value of ( n + 1 / 2 ) th Observations

   = value of ( 5 + 1 / 2 ) th Observations

  = value of 3rd Observations

= 3

Median = 3

c ) Sum of squared deviation

SS = ( \sumx2 -( \sumx)2 / n )

= 71 - ((9 )2 / 5 )

= 71 - (81 / 5 )

= 71 - 16.2

= 54.8

Sum of squared deviation = 54.8

d ) Variance

S2 =   ( \sumx2 -( \sumx)2 / n )n -1 )

= 71 - ((9 )2 / 5 ) 4

= 71 - (81 / 5 )4

= 71 - 16.2 / 4

= 54.8 /4

= 13.7

Variance = 13.7

f ) Standard deviation

S =  \sqrt{} ( \sumx2 -( \sumx)2 / n )n -1 )

= \sqrt{} 71 - ((9 )2 / 5 ) 4

=\sqrt{} 71 - (81 / 5 )4

= \sqrt{} 71 - 16.2 / 4

= \sqrt{} 54.8 /4

= \sqrt{} 13.7

Standard deviation = 3.7014